# Topological Structure Learning Should Be A Research Priority for LLM-Based Multi-Agent Systems

### v1

@@ Appendix A Notations
Table 1 shows the notations used thoughout the paper.
@@
$M$ Macro-structure / coordination pattern
@@
$\gamma$ Cost–reward trade-off coefficient ($\gamma\geq 0$)
$\mathcal{R}(x,M,\hat{\mathcal{A}})$ Expected reward under $M$ with $\hat{\mathcal{A}}$
$\mathcal{U}(x,M,\hat{\mathcal{A}})$ $\mathcal{R}-\gamma\sum C(a_{i})$ utility
$\mu(\hat{\mathcal{A}})$ True expected utility of $\hat{\mathcal{A}}$
@@ Appendix B Algorithm
2:$\hat{\mathcal{A}}\leftarrow\emptyset$, ; $b\leftarrow 0$triangleright $b$ tracks accumulated cost
@@
4:     $\hat{a}\leftarrow\underset{a_{i}\in\mathcal{A}\setminus\hat{\mathcal{A}}}{\mathrm{argmax}}\frac{f(\hat{\mathcal{A}}\cup\{a_{i}\})-f(\hat{\mathcal{A}})}{C(a_{i})}$
5:     if $b+C(a_{i})\leq B$ and $f(\hat{\mathcal{A}}\cup\{a_{i}\})>f(\hat{\mathcal{A}})$ then
6:         $\hat{\mathcal{A}}\leftarrow\hat{\mathcal{A}}\cup\{a_{i}\}$;  $b\leftarrow b+C(a_{i})$
@@
11:return $\hat{\mathcal{A}}$triangleright $(1-1/e)/2$ approximation [40]
@@ C.1 Static Agent Selection
Once cost-constrained exploration stabilizes empirical rewards $\hat{\mu}_{i}$, the problem reduces to choosing a subset $\hat{\mathcal{A}}\subseteq\mathcal{A}$ that maximizes an ensemble utility $\mathcal{U}(\mathcal{A},x)$ subject to the same budget. An exhaustive search over all $\smash{2^{|\mathcal{A}|}}$ subsets is generally intractable [41], but we can exploit the structure of $\mathcal{U}$. Diminishing-returns. In ensembling [42], the marginal benefit of adding one additional agent can diminish due to overlapping knowledge or capabilities. Consequently, $\mathcal{U}(x,M,\hat{\mathcal{A}})$ is approximately submodular in $\hat{\mathcal{A}}$–the benefit of adding an agent decreases as the committee grows. When this holds, the classic greedy algorithm—adding agents one at a time based on their marginal gain–achieves a $(1-1/e)$ approximation to the optimal selection.
@@ C.2 Dynamic Agent Selection
Node Embedding Curation. We propose constructing node embeddings $\mathbf{v}_{i}^{0}$ from four feature groups: 1) Capability: context length, training corpus size, parameter counts, embedding dimensions; 2) Cost: price per 1000 tokens, hardware costs; 3) Performance history: success rate, latency, recent tasks metrics such as F1/BLEU scores [43]; and 4) Persona attributes: communication styles, creativity, safety stance.
Task Bank. Tasks $\mathcal{X}$ are organized in a coarse-to-fine faceted taxonomy: 1) skill: reasoning, code generation, tool use; 2) modality: text, code, image, audio, and video; 3) interaction pattern: single/multi-turn conversations or planning. Each task receives multiple labels across these axes, with associated seed prompts and evaluation rubrics.
Systematic Persona Synthesis. Current approaches to agent selection often rely on handcrafted prompt templates to define personas [44]. For example, creative agents favor broader exploration, while conservative agents prefer safe, high-confidence regions [45]. Agents might prefer clarity (executive), critical thinking (Socratic), or story-telling (narrative).
@@
In both approaches, automated validation–through competency, bias, and efficiency checks–ensures that newly synthesized agents are both safe and effective before deployment [46].
@@
Online Identification via Cost-aware Bandits. Each candidate agent $a_{i}$ is modeled as a stochastic bandit arm with cost $C(a_{i})$ and reward from a utility scoring function. Our objective is to identify a subset of agents that maximizes utility per unit cost across a stream of tasks. Variants based on existing algorithms, such as contextual bandits [47] or Neural UCB [48], guide exploration using query features (e.g., domain tag, length, etc.), prioritizing under-explored yet promising agents. Each round, the controller predicts the expected reward, applies an optimism bonus, and selects the top-$k$ arms. While feedback-graph bandits [49, 50, 51] considered structural interactions among arms, they mostly focus on leveraging this topology to improve sample efficiency by assuming fixed and given graph structures rather than optimizing them. Under a spend budget $B$, budget-aware variants based on budgeted-UCB [52] and budgeted Thompson Sampling [53] manage agent selection by treating each pull as consuming cost $C(a_{i})$ until budget $B$ is reached.
Intermediary Rewards via Automatic Scorers. Human or gold-standard labels offer accurate utility assessments but are costly and sparse. Instead, automatic scorers can be used to provide instantaneous, low-cost reward estimates. Scorer models are architecturally agnostic and can take various forms: a smaller, instruction-tuned specialist model [54, 55], a rule-based evaluator (e.g. exact string match, BLEU/ROUGE, BERTScore, toxicity-detectors), or LLM-as-a-Judge [56]. Critically, the scorer does not need to surpass production agents in size or capability–it merely serves as a lightweight, cost-effective proxy for reward signals. As LLM-based automatic feedback might not always be accurate, we can develop superior mechanisms for providing corrective feedback or rewards to shape agent behaviors in real time without blocking the training flow.
@@
  • Time-varying Connections: Communication patterns between agents can change rapidly (e.g., requests for help, forwarding solutions, ephemeral collaborations). An agent may broadcast to many peers during peak load but fall back to a few trusted collaborators when quotas tighten. TGNNs explicitly encode such time-varying edges.
  • Sequential Dependencies: The impact of a message depends on when it arrives: an expensive agent triggered late in the interaction window may violate a latency SLA even if its answer is accurate. TGNNs condition updates on inter-event time $\Delta t$.
  • Path Dependencies: Agents accumulate context over rounds of communication. TGNNs enable long-range modeling by preserving this information through their node memories.
@@ D.1 Existing Studies
Automated design of agentic systems is a nascent paradigm aiming to replace static, hand-crafted agent organizations with learned ones [55]. AFlow [29] reformulates an agent workflow as a programmatic graph and uses Monte Carlo tree search to iteratively refine this workflow. Similarly, MaAS [23] can sample a query-dependent agent configuration from this supernet at runtime. MasRouter [16] uses a cascaded controller network to dynamically decide how to route a user’s query through a bunch of agents and determine the collaboration mode. G-Designer [26] leverages a graph neural network to generate a communication topology for the agents based on the task at hand. Similarly, CommFormer [27] uses a continuous relaxation of the adjacency matrix to allow gradient-based search of the optimal inter-agent communication links. To guide the formation of an effective communication structure, AOP [24] emphasizes structural principles like solvability, non-redundancy and completeness, where meta agent decomposes and assigns agents. The reinforced meta-thinking agents framework ReMA [25] has a high-level agent that plans and monitors the reasoning strategy, as well as a low-level agent that executes the reasoning steps. The existing work motivates developing a general method to determine optimal multi-agent structures adaptively.
D.2 SPAN
Formulation. We formalize structure profiling as an optimization problem. Let $\mathcal{A}=\{a_{1},a_{2},\dots,a_{N}\}$ be a set of available LLM agents (with possibly diverse expertise or roles), and let $x$ denote an input problem or query. A structure $\mathcal{G}$ describes the organization of these agents for solving $x$, named SPAN (Structure Profiling Adaptive Network). Similar to the temporal graph defined in Section C.2, we can represent $\mathcal{G}$ as a directed graph (or network) $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}\subseteq\mathcal{A}$ is the subset of agents utilized and $\mathcal{E}$ is the set of directed communication links between agents. For example, an edge $e_{ij}\in\mathcal{E}$ indicates agent $a_{i}$’s output is forwarded as input context to agent $a_{j}$, namely $a_{i}\to a_{j}$. This graph can encode various topologies (chain, star, fully connected, etc.) and implicit schedules of agent interactions (a directed acyclic graph imposes a partial order of information flow). Given a particular structure $\mathcal{G}$ and task $x$, the agents execute a collaborative reasoning process (e.g. passing messages along $\mathcal{E}$ and performing computations at each node $a_{i}\in\mathcal{V}$) to produce a final result $y=f_{\theta}(\mathcal{G},x)$.
As the utility function $\mathcal{U}(\mathcal{G};x)$ defined in Section 3.1, it measures the quality of the result (and possibly the efficiency) when using structure $\mathcal{G}$ on task $x$. This could be, for instance, the task reward or accuracy achieved by the agents minus a penalty for communication cost (number of messages or tokens exchanged). The goal of structure profiling is to find the optimal agent network $\mathcal{G}$ for each task. Formally, for a given query $x$ we seek:
$\mathcal{G}^{*}(x)\;=\;\arg\max_{\mathcal{G}\in\mathcal{S}}\;\mathcal{U}(\mathcal{G};x),$ (9)
Let $\mathcal{S}$ be the space of all feasible structures (all possible directed graphs over the agent set $\mathcal{A}$ that respect any required constraints, such as acyclicity or communication budget limits). In words, $\mathcal{G}^{*}(x)$ maximizes the chosen utility, achieving the best trade-off between solution quality and costs for that particular problem. In practice, directly searching $\mathcal{S}$ for each new query is intractable, where $\mathcal{S}$ grows combinatorially with the number of agents and possible connections. Instead, our aim is to learn a structure profiling policy that can rapidly select or construct a high-performing $\mathcal{G}$ for any given task, without brute-force search.
Task-Conditioned Structure Policy. To realize this, we propose a policy function $\Pi$ that maps a task specification to a structured multi-agent configuration. Concretely, $\Pi$ takes as input some representation of the query $x$ (for example, an encoding of the user’s request or a brief summary of the problem) and outputs a definition of the graph $\mathcal{G}$ (which agents to use and how to connect them). We denote this $\mathcal{G}_{\Pi}(x)=\Pi(x)$. The policy can be understood as a planner or graph constructor specialized for assembling LLM agents. In our approach, $\Pi$ is implemented as a parametric model that is trained to approximate the optimal mapping $x\mapsto\mathcal{G}^{*}(x)$.
A convenient formulation is to model $\Pi$ as a function that outputs the probability of forming a communication link between each pair of agents, conditioned on the task. Specifically, for any two agents $a_{i}$ and $a_{j}$, $\Pi$ produces a probability $p_{ij}(x)$ indicating the likelihood that $a_{i}$ should transmit information to $a_{j}$ given task input $x$. Collectively, these probabilities form a task-dependent probabilistic adjacency matrix $P(x)\in[0,1]^{N\times N}$, with the $(i,j)$-th entry of $P(x)$ being $p_{ij}(x)$. A discrete graph $\mathcal{G}$ can then be instantiated by sampling each edge $e_{ij}$ independently with probability $p_{ij}(x)$, or selected deterministically, e.g., including $e_{ij}$ only if $p_{ij}(x)$ exceeds a predefined threshold. In this way, $\Pi$ defines a distribution over topological structures, tailored to each task.
We incorporate domain knowledge and constraints into $\Pi$ to reduce the search space. For instance, if we know certain agents are relevant only for specific domains or sub-tasks, $\Pi$’s architecture can be designed to consider those aspects of $x$. One might use a two-step decision: first choose a subset of agents $\mathcal{V}\subseteq\mathcal{A}$ to be utilized, then decide the directed connections $\mathcal{E}$ among them. This could be achieved by outputting additional binary variables $q_{i}(x)$ indicating whether to activate agent $a_{i}$ for task $x$. The resulting structure policy could be factorized as $\Pi(x)=(\mathcal{V}(x),\mathcal{E}(x))$. For example, $\Pi$ might select agents with the required skills for $x$ (using $q_{i}$ predictions), and then configure a communication graph $\mathcal{E}$ among those agents to best facilitate information flow for the task. We note that many structure profiling strategies can be encoded by an appropriate $\Pi$: a trivial $\Pi$ that always returns a fixed $\mathcal{G}_{0}$ corresponds to a static topology baseline, whereas an ideal $\Pi$ would tailor the graph in a highly task-specific way (e.g. forming teams of experts and a coordinator for a complex analytical question, but choosing a single-agent or linear chain for a simple narrative generation task).
The $\Pi$ might be realized by a graph neural network that, given an embedding of the task (and possibly initial agent state representations), predicts edge scores for each pair of agents [26]. Another approach is to use the LLM agents themselves in a meta-capacity, i.e., one agent (or the system prompt) could be designated to plan the collaboration (writing pseudocode or instructions that effectively set $\mathcal{E}$ and $\mathcal{V}$). While the implementation can vary, the core idea is that $\Pi$ provides a mapping from task features to multi-agent structure in a principled, learnable way.
Optimization Strategy. The parameters of $\Pi$ are trained to maximize the expected utility of the chosen structures. We define an objective across the distribution of tasks (or a training set of tasks $X$):
$J(\Pi)\;=\;\mathbb{E}_{x\sim\mathcal{D}}\big{[}\,\mathcal{U}\!\big{(}\mathcal{G}_{\Pi}(x);\;x\big{)}\,\big{]},$
where $\mathcal{D}$ is the task distribution and $\mathcal{G}_{\Pi}(x)$ is the structure output by policy $\Pi$. Optimizing $J(\Pi)$ aligns the policy to prefer structures that yield high task performance. In practice, this optimization can be approached in different ways depending on how $\Pi$ is implemented. If $\Pi$ is differentiable (e.g., using continuous relaxations for graph decisions [27]), one can apply gradient-based methods to directly increase $\mathcal{U}$. More generally, since $\mathcal{G}$ selection may be discrete, we can employ reinforcement learning that treat the selection of a structure for a given task as an action, receive the utility $\mathcal{U}$ as a reward, and update $\Pi$ to reinforce choices that lead to higher rewards. The policy gradient or evolutionary search can be used to improve $\Pi$ over many sampled tasks. Notably, some recent work has demonstrated the viability of search-based optimization for agent structures (e.g., AFlow’s MCTS optimization of workflows [29]). Our method can leverage similar search procedures: $\Pi$ can initially generate a candidate structure for $x$, which is then refined by local adjustments (adding/removing an edge, activating a new agent, etc.) if those adjustments empirically improve $\mathcal{U}(G;x)$. Such iterative refinement can be guided by heuristics or learned value functions that predict how a structure change would impact future reward.
Importantly, we impose any known constraints and invariance during optimization. For instance, to avoid cycles in communication (which could cause infinite loops or repeated content), we restrict $\mathcal{E}$ to DAGs by design or add a large penalty for cyclic graphs. We may also regularize $\Pi$ to prefer simpler structures (fewer edges) unless additional communication is clearly beneficial, thus preventing gratuitous use of fully connected messaging. The outcome of training is a policy $\Pi^{*}$ that profiles the task structure, i.e., given a new problem, $\Pi^{*}$ can swiftly instantiate a near-optimal multi-agent network for solving it. Formally, we hope to approximate $\mathcal{G}_{\Pi}(x)\approx\mathcal{G}^{*}(x)$ for all $x$ in the domain of interest, without an exhaustive search at query time.
By framing structure profiling as a learning problem, SPAN adapts to the patterns in tasks it has seen. For example, it might learn that analytical math problems warrant a pair of thinker with checker agents, whereas creative writing prompts are best handled by a single agent to maintain coherence. The policy improves its ability to configure the right topology, division of labor, and information routing, which yields a general and extensible solution, that is, as new agent capabilities or tools are added to the pool $\mathcal{A}$, the policy $\Pi$ can learn to incorporate those into structures when beneficial (e.g. including a code-generation agent in the loop only for programming-related queries). The final output of SPAN is therefore not a single static agent system, but a profiling mechanism that flexibly assembles ad hoc agent teams with optimized interaction structures for each task.
D.3 Future Discussion
@@ E.1 Details for Task-Aligned Utility Estimation for Topology Optimization
Counterfactual Topology Intervention. To quantify the topological structure importance of each agent $a_{i}\in\hat{\mathcal{A}}$ or communication edge $e_{i}\in E$ within a synthesized topological structure $G=(\hat{\mathcal{A}},E)\in\mathcal{G}_{M}(\hat{\mathcal{A}})$, we adopt a counterfactual reasoning approach. Specifically, we define the task-aligned contribution of each component, which can be either an agent $a_{i}$ or a communication edge $e$, as the drop in utility resulting from its removal. Let $\mathcal{U}(x,M,\hat{\mathcal{A}})$ denote the utility associated with executing task $x\in T$ under macro-structure M and agent subset $\hat{\mathcal{A}}$. Formally, we define the contribution of each agent $a_{i}\in\hat{\mathcal{A}}$ as
$\phi(a_{i}):=\mathcal{U}(x,M,\hat{\mathcal{A}})-\mathcal{U}(x,M,\hat{\mathcal{A}}\setminus\{a_{i}\}),$ (10)
@@
Here, $\mathcal{U}(G;x)$ denotes the utility under a given micro-topology G, and $G\setminus\{e\}$ or $\hat{\mathcal{A}}\setminus\{a_{i}\}$ respectively represent topologies with the edge or agent removed. Given these contribution scores, we formalize the topology refinement as a constrained selection problem. The goal is to identify a subgraph $G^{\prime}=(\hat{\mathcal{A}}{{}^{\prime}},E{{}^{\prime}})\subseteq G$ that retains only components with sufficiently high task-aligned utility, while maximizing the overall system performance. Let $\delta_{v}$ and $\delta_{e}$ denote contribution thresholds for nodes and edges, respectively. Then, the resulting substructure is obtained by solving:
$G^{*}=\arg\max_{\hat{\mathcal{A}}{{}^{\prime}},E{{}^{\prime}}}\,\mathcal{U}(G^{\prime};x),\>\text{subject to}\;\,\phi(a_{i})\geq\delta_{v},\forall\,a_{i}\in\hat{\mathcal{A}}{{}^{\prime}};\;\,\phi(e)\geq\delta_{e},\forall\,e_{i}\in E{{}^{\prime}}.$ (12)

### v3

@@ Appendix A Notations
Table˜1 shows the notations used thoughout the paper.
@@
$M$ Macro‑structure / coordination pattern
@@
$\gamma$ Cost–reward trade‑off coefficient ($\gamma\geq 0$)
[…]
@@ Appendix B Algorithm
[…]
@@
[…]
@@
11:return $\hat{\mathcal{A}} \triangleright$ (11/e)/2(1-1/e)/2 approximation (Sviridenko, 2004)
@@ C.1 Static Agent Selection
[…]
@@ C.2 Dynamic Agent Selection
Node Embedding Curation. We propose constructing node embeddings $\mathbf{v}_{i}^{0}$ from four feature groups: 1) Capability: context length, training corpus size, parameter counts, embedding dimensions; 2) Cost: price per 1000 tokens, hardware costs; 3) Performance history: success rate, latency, recent tasks metrics such as F1/BLEU scores (Papineni et al., 2002); and 4) Persona attributes: communication styles, creativity, safety stance.
Task Bank. Tasks $\mathcal{X}$ are organized in a coarse‐to‐fine faceted taxonomy: 1) skill: reasoning, code generation, tool use; 2) modality: text, code, image, audio, and video; 3) interaction pattern: single/multi-turn conversations or planning. Each task receives multiple labels across these axes, with associated seed prompts and evaluation rubrics.
Systematic Persona Synthesis. Current approaches to agent selection often rely on handcrafted prompt templates to define personas (Jin et al., 2024b). For example, creative agents favor broader exploration, while conservative agents prefer safe, high-confidence regions (Sun et al., 2024). Agents might prefer clarity (executive), critical thinking (Socratic), or story-telling (narrative).
@@
In both approaches, automated validation–through competency, bias, and efficiency checks–ensures that newly synthesized agents are both safe and effective before deployment (Jin et al., 2024a).
@@
Online Identification via Cost-aware Bandits. Each candidate agent $a_{i}$ is modeled as a stochastic bandit arm with cost $C(a_{i})$ and reward from a utility scoring function. Our objective is to identify a subset of agents that maximizes utility per unit cost across a stream of tasks. Variants based on existing algorithms, such as contextual bandits (Chu et al., 2011) or Neural UCB (Zhou et al., 2020), guide exploration using query features (e.g., domain tag, length, etc.), prioritizing under-explored yet promising agents. Each round, the controller predicts the expected reward, applies an optimism bonus, and selects the top-$k$ arms. While feedback-graph bandits (Alon et al., 2015; Cesa-Bianchi et al., 2021; Rouyer et al., 2022) considered structural interactions among arms, they mostly focus on leveraging this topology to improve sample efficiency by assuming fixed and given graph structures rather than optimizing them. Under a spend budget $B$, budget-aware variants based on budgeted-UCB (Ding et al., 2013) and budgeted Thompson Sampling (Xia et al., 2015) manage agent selection by treating each pull as consuming cost $C(a_{i})$ until budget $B$ is reached.
Intermediary Rewards via Automatic Scorers. Human or gold-standard labels offer accurate utility assessments but are costly and sparse. Instead, automatic scorers can be used to provide instantaneous, low-cost reward estimates. Scorer models are architecturally agnostic and can take various forms: a smaller, instruction-tuned specialist model (Li et al., 2025; Hu et al., 2024b), a rule-based evaluator (e.g. exact string match, BLEU/ROUGE, BERTScore, toxicity-detectors), or LLM-as-a-Judge (Zheng et al., 2023). Critically, the scorer does not need to surpass production agents in size or capability–it merely serves as a lightweight, cost-effective proxy for reward signals. As LLM-based automatic feedback might not always be accurate, we can develop superior mechanisms for providing corrective feedback or rewards to shape agent behaviors in real time without blocking the training flow.
@@
  • Time-varying Connections: Communication patterns between agents can change rapidly (e.g., requests for help, forwarding solutions, ephemeral collaborations). An agent may broadcast to many peers during peak load but fall back to a few trusted collaborators when quotas tighten. TGNNs explicitly encode such time‐varying edges.
  • Sequential Dependencies: The impact of a message depends on when it arrives: an expensive agent triggered late in the interaction window may violate a latency SLA even if its answer is accurate. TGNNs condition updates on inter‑event time $\Delta t$.
  • Path Dependencies: Agents accumulate context over rounds of communication. TGNNs enable long‐range modeling by preserving this information through their node memories.
@@ D.1 Existing Studies
Automated design of agentic systems is a nascent paradigm aiming to replace static, hand-crafted agent organizations with learned ones (Hu et al., 2024b). AFlow (Zhang et al., 2024c) reformulates an agent workflow as a programmatic graph and uses Monte Carlo tree search to iteratively refine this workflow. Similarly, MaAS (Zhang et al., 2025) can sample a query-dependent agent configuration from this supernet at runtime. MasRouter (Yue et al., 2025) uses a cascaded controller network to dynamically decide how to route a user’s query through a bunch of agents and determine the collaboration mode. G-Designer (Zhang et al., 2024b) leverages a graph neural network to generate a communication topology for the agents based on the task at hand. Similarly, CommFormer (Hu et al., 2024a) uses a continuous relaxation of the adjacency matrix to allow gradient-based search of the optimal inter-agent communication links. To guide the formation of an effective communication structure, AOP (Li et al., 2024) emphasizes structural principles like solvability, non-redundancy and completeness, where meta agent decomposes and assigns agents. The reinforced meta-thinking agents framework ReMA (Wan et al., 2025) has a high-level agent that plans and monitors the reasoning strategy, as well as a low-level agent that executes the reasoning steps. The existing work motivates developing a general method to determine optimal multi-agent structures adaptively.
D.2 Future Discussion
@@ E.1 Details for Task-Aligned Utility Estimation for Topology Optimization
[…]
@@
[…]
$G^{*}=\arg\max_{\hat{\mathcal{A}}{{}^{\prime}},E{{}^{\prime}}}\,\mathcal{U}(G^{\prime};x),\>\text{subject to}$ (8)
$\;\,\phi(a_{i})\geq\delta_{v},\forall\,a_{i}\in\hat{\mathcal{A}}{{}^{\prime}};\;\,\phi(e)\geq\delta_{e},\forall\,e_{i}\in E{{}^{\prime}}.$
@@ E.2 Future Discussion
Appendix F Use of Large Language Models
We used ChatGPT solely to refine the manuscript’s language–improving grammar, phrasing, and clarity. All technical content, experimental design, analyses, and conceptual contributions were authored and verified by the authors.